A car is accelerating in the +x direction. At 2s the car is 100m from the starting point. After an additional 5s, it is 1.5km from the starting point.






Determine the average velocity of the car for the first 5s of its travel.






Determine the average velocity of the car for the entire period of its travel.

1
Expert's answer
2022-01-26T12:17:01-0500

"V_0=\frac{x_0}{t}= \frac{100}{2} = 50 \frac{m}{sec^2}"


The average velocity of the car for the first 5s of its travel.


"V_1=\frac{x_1}{t_1}= \frac{1500}{5} = 300 \frac{m}{sec^2}"


 The average velocity of the car for the entire period of its travel.


"V= \frac{V_0+V_1}{2}=\frac{350}{2} = 175 \frac{m}{sec^2}"